# ImageNot: A contrast with ImageNet preserves model rankings

### v1

@@ A.2 Trained Models
Table 2: Hyperparameters search space and Training Budgets for experiment. TL: Transfer Learning; FT: Finetuning; ST: Standard (train/test with randomly initialized weights).
@@
ImageNot batch size ST [256] 4
ImageNot optimizer ST [SGD] 4
ImageNot lr ST [5e-1, 1e-1, 1e-2, 1e-4] 4
ImageNot lr decay ST [0.1] 4
ImageNot lr scheduler ST [Linear Warm Start + Cosine Annealing] 4
ImageNot momentum ST [0.9] 4
ImageNot optimer warmup steps ST [5] 4
ImageNot weight decay ST [2e-5] 4
(*) For the reported batch size in Table 2, we use [128/32/32] for EfficientNet-V2-L and [256/64/64] for ConvNeXt-large.
We utilize early stopping with a patience of 10 on the basis of held-out validation sample accuracy and set a maximum epoch of 1000 for all training. Additionally, to optimize training efficiency and reduce computational costs, we employ the Asynchronous Hyper-Band Scheduler from Ray Tune (Li et al., 2020; Liaw et al., 2018), which enables us to further early-stop less promising hyperparameters.
Appendix B Additional Results and Discussion
@@
Table 3 summarizes the differences (or lack there of) between properties of the ImageNet and ImageNot WordNet synsets (referenced in section 3.1).
B.2 Robustness
Robustness Ranking.
We employ 17 distinct test-time data augmentations to independently rank model accuracies, all random except for the Identity augmentation. These augmentations are applied to the same set of held-out test examples. They include: (1) Identity, which uses the original image; (2) Gaussian Blur, applying a Gaussian blur to smooth the image and reduce noise; (3) Adjust Sharpness, for modifying image sharpness; (4) Affine, performing affine transformations such as scaling, rotations, and translations; (5) ColorJitter, adjusting brightness, contrast, saturation, and hue; (6) Contrast, altering contrast levels; (7) Equalize, for histogram adjustment and contrast balancing; (8) Grayscale, converting the image to grayscale; (9) Horizontal Flip, flipping the image horizontally; (10) Invert, inverting pixel values; (11) Perspective, applying perspective transformations; (12) Posterize, reducing the bit depth per color channel; (13) Resized Crop, cropping after resizing; (14) Rotation, rotating the image; (15) Sharpness, for sharpness adjustment; (16) Solarize, inverting colors post a threshold; and (17) Vertical Flip, flipping the image vertically.
We take the average across these augmentations to generate our robustness rankings. The model rankings are surprisingly stable on average across these 17 test-time augmentations (Figure 10-12).
Figure 10: Models initialized with random weights. Train on ImageNet. Evaluate on ImageNet test set.
We find that models trained on ImageNet maintain rankings on average when evaluated on the ImageNet test sets with test-time data corruption (Figure 10).
Figure 11: Models initialized with random weights. Train on ImageNot. Evaluate on ImageNot test set.
Additionally, we find that models trained on ImageNot maintain rankings on average when evaluated on the ImageNot test sets with test-time data corruption (Figure 11). However, across the corruptions, the rankings are less stable than what we observed for ImageNet. We hypothesize that this is a consequence of our guardrails on model tuning, which leads to sub-optimal models than those that could have been achieved with more extensive tuning.
Figure 12: Models initialized with random weights. Train on CIFAR10. Evaluate on CIFAR10 test set.
Finally, we find that models trained on CIFAR10 maintain rankings on average when evaluated on the CIFAR10 test sets with test-time data corruption (Figure 12). However, across the corruptions, the rankings are less stable than what we observed for ImageNet. We hypothesize that this is a consequence of our guardrails on model tuning, which leads to sub-optimal models than those that could have been achieved with more extensive tuning.

### v2

@@ A.2 CIFAR10 Models
Table 2: Hyperparameters search space and Training Budgets for experiment. TL: Transfer Learning to CIFAR10; FT: Finetuning to CIFAR10; ST: Standard, randomly initialized weights trained only on CIFAR10.
@@
For both fine-tuning and transfer learning from ImageN[o/e]t, we use these pretrained weights as our starting point and then update weights either by fine-tuning or transfer learning. We also train end-to-end only on CIFAR10 with randomly initialized model weights. The hyperparameter details can be found in Table 2.
A.3 Robustness
We employ 17 distinct test-time data augmentations to evaluate the robustness of the developed models, all random except for the Identity augmentation. These augmentations are applied to the same set of held-out test examples. They include: (1) Identity, which uses the original image; (2) Gaussian Blur, applying a Gaussian blur to smooth the image and reduce noise; (3) Adjust Sharpness, for modifying image sharpness; (4) Affine, performing affine transformations such as scaling, rotations, and translations; (5) ColorJitter, adjusting brightness, contrast, saturation, and hue; (6) Contrast, altering contrast levels; (7) Equalize, for histogram adjustment and contrast balancing; (8) Grayscale, converting the image to grayscale; (9) Horizontal Flip, flipping the image horizontally; (10) Invert, inverting pixel values; (11) Perspective, applying perspective transformations; (12) Posterize, reducing the bit depth per color channel; (13) Resized Crop, cropping after resizing; (14) Rotation, rotating the image; (15) Sharpness, for sharpness adjustment; (16) Solarize, inverting colors post a threshold; and (17) Vertical Flip, flipping the image vertically.
Appendix B Additional Results
Figure 9: Final Model Ranking and Relative Improvement. Model rankings and relative improvements hold with models trained from scratch on the respective dataset, and evaluated on a held-out test set. For a set of test accuracies $X$, relative progress is $X_{i}/X_{\text{AlexNet}}$. ImageNet model accuracies range from 0.57 (AlexNet) to 0.85 (EfficientNet V2 L), while much noisier ImageNot model accuracies approximately range from 0.4 (AlexNet) to 0.6 (EfficientNet V2 L). CIFAR10 accuracies range from 0.89 (AlexNet) to 0.95 (ConvNeXt Large). The rankings for CIFAR10 are derived from means across three restarts; we note that ranking differences from ImageNet only occur for models with error bounds overlapping with the model in its ImageNet position. Thus, even for CIFAR10, the rankings are preserved up to error. Robust corresponds to average accuracy across many test-time distribution shifts.
@@ Appendix B Additional Results
Table 3 summarizes the differences (or lack thereof) between properties of the ImageNet and ImageNot WordNet synsets (referenced in section 3.1). Figures 11 and 12 further depict the difference in ImageNot and ImageNet’s WordNet subtree structure.
Figure 11: Synset depth in WordNet tree.
Figure 12: Number of synset siblings.